OpenStudy (anonymous):

What are the zeros of the polynomial function: f(x) = x3 – x2 – 12x ?

OpenStudy (anonymous):

i found the answer but thanks

OpenStudy (anonymous):

First, we see that (x) exist in all the terms, so we can factor out an (x) and get: f(x) = x^3 - x^2 - 12x = x( x^2 - x -12) Now we factor apart the (x^2 - x -12). What two numbers added together give -1 and multiplied together gives -12? I think it is -4 and 3, and I am pretty sure you go that too. so: x(x-4)(x+3), which means x = 0, 4,-3